3. Support Department Cost Allocation—Reciprocal Services Method

Brewster Toymakers Inc. produces toys for children. The toys are produced in the Molding and Assembly departments. The Janitorial and Security departments support the production of the toys. Costs from the Janitorial Department are allocated based on square feet. Costs from the Security Department are allocated based on asset value. Relevant department information is provided in the following table:

Janitorial
Department
Security
Department
Molding
Department
Assembly
Department
Square feet 600    3,600    3,600    1,800   
Asset value $100    $120    $900    $1,000   
Department cost $2,100    $1,800    $10,400    $12,700   

Using the reciprocal services method of support department cost allocation, determine:

a. The percentage of Janitorial costs that should be allocated to the Security Department.

b. The percentage of Security costs that should be allocated to the Janitorial Department.

4. Joint Cost Allocation—Physical Units Method

Blake’s Blacksmith Co. produces two types of shotguns, a 12-gauge and 20-gauge. The shotguns are made through a joint production process that ultimately produces 35 12-gauge shotguns and 35 20-gauge shotguns and costs a total of $3,700 per batch. After the split-off point, each type of shotgun goes through an additional crafting process before it is sold. The additional production process of the 12-gauge shotgun costs $35 per gun, after which it is sold for $184 per gun. The additional production process of the 20-gauge shotgun costs $24 per gun, after which it is sold for $154 per gun.

Determine the amount of joint production costs allocated to each type of shotgun using the physical units method.

Joint Product Allocation
12-gauge shotgun $     
20-gauge shotgun      
Totals $     

5. Joint Cost Allocation—Weighted Average Method

Gary’s Grooves Co. produces two types of carving knives, one with a handle made of a polymer that looks like walnut wood and another with a handle made with a polymer that looks like red oak. The knives are made through a joint production molding process that produces 301 knife blades for red oak handle knives and 258 knife blades for walnut handle knives at the split-off point. The polymer for the red oak handle knife blades requires twice as much cooling time as the polymer for the walnut handle knife blades, although all knives are removed from the joint molding process at the same time (i.e., once the cooling for the red oak handle knives is complete). The joint production process costs a total of $5,000.

Assuming the company allocates joint costs using the weighted average method based on the required cooling time of the two joint products, determine the amount of joint production costs allocated to each type of knife using the weighted average method.

Joint Product Allocation
Red oak handle carving knife $
Walnut handle carving knife
Totals $
Answer #1

3.

a.

Percentage of Janitorial costs to be allocated to Security Department = 3,600 / (3,600 + 3,600 + 1,800) = 40%

b.

Percentage of Security costs to be allocated to Janitorial Department = 100 / (100 + 900 + 1,000) = 5%

.

4.

Joint cost allocation
Physical units Joint cost allocated
12-gauge 35 $1,850 ($3,700 * 35 / 70)
20-gauge 35 $1,850 ($3,700 * 35 / 70)
Total 70 $3,700

.

5.

Gary’s Grooves Red Oak Walnut Total Note
Units Produced      301.00       258.00 A
Weights assigned           2.00           1.00 B
Weighted production units      602.00       258.00       860.00 C=A*B
Joint costs 5,000.00 D
Cost per Weighted units           5.81 E=D/C
Cost per units         11.63           5.81 F=E*B
Cost allocated       3,500        1,500 5,000 G=F*A

.

*** End ***

Do leave a thumbs up :)